OpenStudy (anonymous):

please help! medals rewarded! The ratio of the measures of two supplementary angles is 7:3. Find the measures of the angles. A. 126° and 54° B. 72° and 18° C. 159° and 21° D. 69° and 21°

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

7x+3y=180

OpenStudy (anonymous):

wich pair works when you put in the equation

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

i am just checking if 72:18 is 7:3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

supplementary totals 180 - so you can ignore B and D

OpenStudy (anonymous):

it isnt so check another

OpenStudy (anonymous):

what is 126/180 and 159/180 = correct one is .70

OpenStudy (anonymous):

So what choice would it be?

OpenStudy (anonymous):

ohh wait correct formula \[7k+3k=180\] \[10k=180\]\ \[k=18\] so the angles are\[7*18\] and \[18*3\]

OpenStudy (anonymous):

so A then!:) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

dont give medals we dont need

OpenStudy (anonymous):

thanks!:)

OpenStudy (anonymous):

so for this question The ratio of the measures of two complementary angles is 1:5. What are the measures of the angles? A. 30° and 60° B. 15° and 75° C. 30° and 150° D. 5° and 25° I got the answer B is that correct? @Jonask

OpenStudy (anonymous):

now its complementary so 90 k+5k=90 6k=90 k=15 1*15 and 5*15

OpenStudy (anonymous):

yes correct